# Arrow-Debreu Meets Kyle: Price Discovery Across Derivatives

### 06 Mar 2026 v41

@@ 8 Empirical Implications
The intuition is straightforward. High expected volatility makes a straddle-type payoff especially valuable, so the insider concentrates demand in the wings by buying deep in-the-money (ITM) and out-of-the-money (OTM) options. This pushes up wing option prices and therefore their implied volatilities. Cross-strike price impact then propagates this demand across the surface: because payoffs at deep ITM and deep OTM strikes are positively correlated (as illustrated by strikes $x_{1}$ and $x_{3}$ in Figure 4), order flow in one wing raises prices in the other wing as well. Importantly, these cross-strike spillovers become stronger farther from at-the-money (ATM), since payoff covariance increases with distance from ATM. The result is the characteristic convexity of the implied-volatility curve—a smile produced by information aggregation across strikes rather than by strike-by-strike pricing in isolation.
A further implication is a sharp comparative static for smile dynamics. As true volatility rises, the implied-volatility curve becomes more convex (see Figure 7b). [^30] As with our other empirical hypotheses, this prediction can be taken directly to the data. Fix a maturity $T$. For each day $t$, construct a measure of smile curvature $\kappa_{t}(T)$, e.g., $\kappa_{t}(T)=\tfrac{1}{2}\big(\sigma^{IV}_{t}(-\Delta,T)+\sigma^{IV}_{t}(+\Delta,T)\big)-\sigma^{IV}_{t}(0,T)$ for a fixed delta $\Delta$ (or the quadratic curvature coefficient from a regression of implied volatility on log-moneyness). Estimate $\widehat{\Lambda}_{t}(T)$ from (38) using intraday option midquote changes and signed order imbalance, and form a cross-strike transmission statistic $\widehat{CI}^{\text{wing}}_{t}(T)$ (e.g., an average of symmetrized cross impacts between the two wings, or between each wing and ATM). The model predicts that curvature co-moves positively with $\widehat{CI}^{\text{wing}}_{t}(T)$ and steepens in high-volatility states, which can be tested in regressions such as $\kappa_{t}(T)=\alpha+\beta\,\widehat{CI}^{\text{wing}}_{t}(T)+\delta\,\text{Vol}_{t}+\eta\big(\widehat{CI}^{\text{wing}}_{t}(T)\times\text{VolState}_{t}\big)+\Gamma^{\prime}X_{t}+\varepsilon_{t},$ where $\text{Vol}_{t}$ is realized volatility and $X_{t}$ includes standard liquidity controls (spreads, depth, volume, etc.). The intuition is immediate: higher volatility increases the covariance between payoffs at deep ITM and deep OTM strikes, which strengthens cross-strike price impact via the propagation of informed demand across strikes. In this sense, changes in smile curvature are the observable footprint of information aggregation across option markets—rather than a reduced-form feature to be imposed ex ante.
@@ 9 Conclusion
The tractability and generality of our model invite natural follow-on questions. How does risk aversion reshape informed portfolios and the pricing kernel, and how does asymmetric information distort risk sharing across contingent claims? In a dynamic setting, how do information spillovers propagate across strikes and maturities? Beyond equity options, the same approach can be applied to futures and forwards, credit derivatives, and interest-rate derivatives, where order-flow-driven prices map into term-structure, default-intensity, and correlation information. We hope to pursue these and related questions in future work.

### 20 Mar 2026 v42

@@ 8 Empirical Implications
The mechanism is straightforward. For a given volatility state, demand shifts toward option positions that load on large price moves, raising option prices and implied volatilities in the put and call wings. Cross-strike price impact then propagates this pressure across the curve: because payoffs at deep in-the-money and deep out-of-the-money strikes are positively correlated (as illustrated by strikes $x_{1}$ and $x_{3}$ in Figure 4), order flow in one wing raises prices in the other wing as well. Moreover, these cross-wing spillovers intensify farther from at-the-money, where payoff covariance is higher. The result is the characteristic convexity of the implied-volatility curve—a smile that emerges from the aggregation of volatility information across strikes.
Across volatility states, the same mechanism implies a clear comparative static: higher volatility leads to a more convex smile (see Figure 7b). The reason is that higher volatility increases the covariance between payoffs at deep in-the-money and deep out-of-the-money strikes, thereby amplifying cross-strike price impact. As a result, smile curvature rises with the intensity of volatility information being aggregated across strikes. Variation in smile curvature is thus the observable footprint of that aggregation. [^30] This comparative static is again testable. Fix a maturity $T$. For each day $t$, estimate $\widehat{\Lambda}_{t}(T)$ from (38) and construct a cross-wing impact measure $\widehat{CI}^{\mathrm{wing}}_{t}(T)$, for example, as the average cross-impact between the left and right wings. The model predicts that smile curvature at maturity $T$ is increasing in $\widehat{CI}^{\mathrm{wing}}_{t}(T)$ and is higher in high-volatility states.
@@ 9 Conclusion
The tractability and generality of our model invite natural follow-on questions. How does risk aversion reshape informed portfolios and the pricing kernel, and how does asymmetric information distort risk sharing across contingent claims? In a dynamic setting, how do information spillovers propagate across strikes and maturities? Beyond equity options, the same approach can be applied to futures and forwards, credit derivatives, and interest-rate derivatives, where order-flow-driven prices map into term-structure, default-intensity, and correlation. We hope to pursue these and related questions in future work.